OpenStudy (anonymous):

The English Alphabet has 20 letters more than one-fourth the number of letters in the Greek Alphabet. The English Alphabet has 26 letters. Which of the following equations would you use to find the number of letters, g, in the Greek Alphabet? Answer 1/4g + 20 = 26 4g - 26 = 20 4g - 20 = 26 g+20/4=26

OpenStudy (lilg132):

1/4g + 20 = 26

hero (hero):

lilg132, i posted a solution to that mixture problem

OpenStudy (lilg132):

which one?

hero (hero):

The one that you started to respond to but didn't

OpenStudy (lilg132):

lol oh that one ye will have a look at it in a minute

OpenStudy (lilg132):

thanks

OpenStudy (anonymous):

so.... is that the answer? or is it a diffrent one?

OpenStudy (anonymous):

?

hero (hero):

checking...

hero (hero):

1/4g + 20 = 26

hero (hero):

should be written as g/4 + 20 = 26

OpenStudy (lilg132):

already answered it again with an explantaion

hero (hero):

She wanted someone else to verify the answer though

hero (hero):

Do you have an issue with that?

OpenStudy (lilg132):

hero do you something up ur a5s everytime i answer a question? or is ur head so far up ur a5s that you get annoyed when someone else answers instead of you she only wanted it verified because you came into this question talking about another question hence why she was so confused now try and pull your head out and carry on try not to get upset

hero (hero):

What? That's not why. Why are you so angry. I think you're overreacting clearly.

hero (hero):

You know, you didn't have to respond after I responded. She simply wanted the answer verified because I came here and she saw that I was here. You are being silly.